OpenStudy (anonymous):

solve and check your answer 2/1-x=5/1+x the solution is x= there is no solution

OpenStudy (anonymous):

2-x = x+5

OpenStudy (anonymous):

2 - x == 5 + x

OpenStudy (anonymous):

first look for restrictions,x#1,x#-1 multiply by (1-x)(1+x) simplify ans solve

OpenStudy (anonymous):

x=1?

OpenStudy (anonymous):

sorry just getting the hang of these

OpenStudy (anonymous):

is it: 2/(1-x)=5/(1+x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok